OpenStudy (anonymous):

\[\sqrt{x-4}-\sqrt[4]{x-4}=6\] why is the solution 85

OpenStudy (anonymous):

please help and list each step, I am so confused

OpenStudy (helder_edwin):

u have \[\large \sqrt{x-4}-\sqrt[4]{x-4}=6 \] which can be rewritten as \[\large \sqrt{x-4}-\sqrt{\sqrt{x-4}}=6 \]

OpenStudy (helder_edwin):

now let \(Y=\sqrt{x-4}\) so u have \[\large Y-\sqrt{Y}=6 \] or \[\large Y-6=\sqrt{Y} \] square both sides \[\large Y^2-12Y+36=Y \] or \[\large Y^2-13Y+36=0 \]

OpenStudy (anonymous):

I still don't understand?

OpenStudy (helder_edwin):

this can be factored into (Y-4)(Y-9)=0

OpenStudy (helder_edwin):

if Y=4 then \(\sqrt{x-4}=4\) gives you x=20

OpenStudy (helder_edwin):

if Y=9 then \(\sqrt{x-4}=9\) gives you x=85

OpenStudy (helder_edwin):

which part don't u understand? (sorry i'am having some problems posting things.)

OpenStudy (anonymous):

oh i understand now!! I didn't realize the factoring